OpenStudy (anonymous):

Find the vertex of the parabola. y=-4x^2-16x-11

myininaya (myininaya):

y=-4(x^2+4x)-11 y=-4(x^2+4x+4)-11+4(4) y=-4(x+2)^2+5 Vertex is (-2,5)

OpenStudy (anonymous):

Thank You thats what I came up with

OpenStudy (anonymous):

Find the slope of the line passing through the points (-3, -10) and (4, 5). (Points : 1)